### v1

# Explainability of Machine Learning Approaches in Forensic Linguistics: A Case Study in Geolinguistic Authorship Profiling

Abstract
Forensic authorship profiling uses linguistic markers to infer characteristics about an author of a text. This task is paralleled in dialect classification, where a prediction is made about the linguistic variety of a text based on the text itself. While there have been significant advances in the last years in variety classification (Jauhiainen et al., 2019) and state-of-the-art approaches reach accuracies of up to 100% depending on the similarity of varieties and the scope of prediction (e.g., Milne et al., 2012; Blodgett et al., 2017), forensic linguistics rarely relies on these approaches due to their lack of transparency (see Nini, 2023), amongst other reasons. In this paper we therefore explore explainability of machine learning approaches considering the forensic context. We focus on variety classification as a means of geolinguistic profiling of unknown texts. For this we work with an approach proposed by Xie et al. (2024) to extract the lexical items most relevant to the variety classifications. We find that the extracted lexical features are indeed representative of their respective varieties and note that the trained models also rely on place names for classifications.
1 Introduction
Forensic authorship analysis, i.e., the analysis of texts to gain information about the author(s) of those texts, is a central domain of research within forensic linguistics. Generally, there are three main areas within authorship analysis working to infer information from questioned documents: Firstly, authorship attribution is interested in learning who the most likely candidate author is given a comparison with the writings of a set of potential authors; Secondly, authorship verification focuses on evaluating if texts were authored by the same person; And thirdly, authorship profiling works to infer characteristics of an author by their language use (Roemling and Grieve, 2024).
While authorship attribution and verification rely on comparisons between (groups of) texts, profiling works only with the questioned document(s) itself. Generally, all areas of authorship analysis can be approached qualitatively and/or quantitatively if the amount of data available permits quantitative analysis. For example, quantitative work in authorship profiling has researched inferring age or gender (Nini, 2018) or native language influence (Perkins and Grant, 2018) from questioned documents.
Nevertheless, forensic authorship profiling is often done in a manual or qualitative way, relying on the expertise of the forensic linguist as can be seen in famous examples like the Unabomber analysis (Leonard et al., 2017) or the devil strip ransom note (Shuy, 2001). In both cases, law enforcement was interested in the regional origin as one of the relevant characteristics of the authors. This background can be inferred through analysing the regional linguistic variation, i.e., the use of regional dialect, in the questioned documents. This is referred to as regional or geolinguistic profiling (Roemling and Grieve, 2024) and is a parallel task to inferring the regional variety of a document as is done in dialect classification or language variety identification (Jauhiainen et al., 2019).
Even though research in forensic linguistics works more and more with statistical and computational approaches (e.g., Bevendorff et al., 2023; Ishihara et al., 2024), authorship profiling often remains a manual task. This is at times credited to the black-box approaches in current NLP research, meaning that the lack of explainability precludes these approaches from being used in legal settings (see Nini, 2023).
2 Related Work
[…]
Xie et al. (2024) were the first to research explainability in a dialect classification context. In order to analyze the classifier, they extracted lexical features that were highly relevant to the classification aiming to use this knowledge for dialect research and not only as a means to understand machine learning approaches. They stated that they relied on lexical items, owing to the complex nature of handling preprocessing like parsing or POS-tagging in low(er)-resourced varieties and indicated that refined approaches would be beneficial. However, previous research on regional variation using social media data has shown that approaches using lexical features provided excellent results (e.g., Doyle, 2014; Huang et al., 2016; Eisenstein, 2017; Grieve et al., 2018, 2019).
[…]
In a forensic authorship profiling setting, an approach like this seems beneficial. It reaches high accuracies in language variety identification, while similarly providing explanations of which features predominantly influenced the classification. The approach also has the advantage of eliminating or at least minimising researcher bias as the models learn the relevant features themselves, whereas it is the forensic linguist’s expertise that culls the features in a qualitative analysis (see Grant and Grieve, 2022) [^1] Note, however, that the training data itself may introduce bias into the model (see Blodgett et al., 2020).. While the approach does not fully explain the inner workings of the model, experts can use the extracted features to a) verify that the model indeed reached a sound decision, for example by evaluating the features against previous dialectological findings, and b) use the explanations to introduce the method to law enforcement or jurisprudence. Even if the classifiers themselves do not meet court admissibility standards (Coulthard, 2013; Hammel, 2022), extracted features can be used for authorship work to contribute to a more objective analysis.
3 Data
We work with a corpus of German social media data, which was collected by Hovy and Purschke (2018); Purschke and Hovy (2019) from the platform Jodel and which has also been used in VarDial classification tasks (Gaman et al., 2020; Chakravarthi et al., 2021). Jodel is structurally similar to Twitter/X, however it only allows anonymous posts. Users of Jodel can interact with other users in a 10-15 km radius around their own location, so all posts are geolocated. The corpus contains posts from Austria, Germany and Switzerland. While most of the data is written in standard German, it shows clear regional patterns. Especially in Switzerland, Austria and, in parts, Bavaria writing is considerably further from standard German (Purschke and Hovy, 2019). Posts from Romandy contain substantial amounts of French.
The corpus consists of approximately 240 million tokens from about 8500 locations, however only 388 locations have a token count of over 10k. For our classification experiments, we mapped these locations onto wider dialect regions following three settings with 3, 4 and 5 classes respectively. The 3-class distinction is based on national borders, so the classes reflect Austria, Germany and Switzerland. In the 4-class setting Germany is additionally divided into two parts, north and south (at latitude 50.33° N) and for the 5-class setting the southern region of Germany is further split into east and west (at longitude 9.97° E). These divisions were operationalized based on knowledge from traditional dialectology (Wiesinger, 1983; König, 2004).
We randomly sampled 200k examples per class for training, and 20k examples per class for development and testing, respectively. Besides some simple whitespace normalization (i.e., removing line breaks and tabs inside a message), we did not perform any preprocessing.
4 Methods and Experiments
4.1 Dialect classifiers
The approach described by Xie et al. (2024) starts by fine-tuning a BERT-based language model on the dialect classification task. Following Xie et al. (2024), we rely on the multilingual model xlm-roberta-base [^2] https://huggingface.co/FacebookAI/xlm-roberta-base (Conneau et al., 2020) but we also experiment with a base model specifically trained on German data, dbmdz/bert-base-german-cased [^3] https://huggingface.co/dbmdz/bert-base-german-cased.
Considering two base models and three settings (3, 4 and 5 classes), our experiments yield six dialect classifiers. The training is done with the simpletransformers library [^4] https://simpletransformers.ai/. Each model is trained for 10 epochs with a maximum sequence length of 256 (subword tokens) and a batch size of 64 samples. We use default values for all other parameters.
Classes Random XLM-RoBERTa German BERT 3 33% 75.31% 74.91% 4 25% 58.57% 58.44% 5 20% 47.74% 47.64%
Table 1: Classification accuracies on the development sets.
Table 1 shows the classification accuracies of these models on the development sets. It can be seen that all models outperform the random baseline by a large margin. The difference between the two base models is marginal and we thus focus on the German base model.
4.2 Leave-one-word-out classification
The LOO method used by Xie et al. (2024) processes each sentence of the test set independently to detect the words that contribute most to the classification. It consists of the following steps:
  1. 1.
    Select an instance $x$ of the test set, run it through the dialect classifier, and record the predicted class $\hat{y}$ as well as the prediction score $\ell$. If the prediction is incorrect ($\hat{y}\neq y$), skip this instance and move to the next one. [^5] This corresponds to the isCorrect constraint of Xie et al. (2024).
  2. 2.
    Select one word of the instance and remove it from the instance (let $x_{i}$ denote the instance $x$ from which word $i$ is removed), run it through the dialect classifier, and record the prediction score $\ell_{i}$. [^6] Xie et al. (2024) select the word to be removed by position, with the consequence that if a word occurs several times in the same sentence, only one of its occurrences will be removed at a time. They then only consider the occurrence that produced the biggest difference. We simplify this part by iterating over the set of unique words and removing all occurrences of the selected word at the same time.
  3. 3.
    Measure the impact of the removed word on the classification performance ($\Delta_{i}$) by subtracting the score of the incomplete instance from the score of the complete instance: $\Delta_{i}=\ell-\ell_{i}$. We call $\Delta_{i}$ the impact score of word $i$.
  4. 4.
    Repeat steps 2 and 3 for each word of the sentence.
  5. 5.
    Select the 5 words with the highest impact score. [^7] Xie et al. (2024) omit this step in the description of their work, but it is present in their code.
The steps described above produce explanations at instance level, i.e., the most impactful words of each instance. This can be an interesting result in itself, but for the analysis reported here, we aggregate the explanations across all instances of the test set. The resulting list is then processed in the following way:
  1. 1.
    Words that were selected as explanations for more than one class are eliminated from further consideration. [^8] This corresponds to the isUnique constraint of Xie et al. (2024).
  2. 2.
    Words that figure as explanations for only one instance are eliminated from further consideration.
  3. 3.
    For each remaining word, we compute the average impact score on the basis of the individual impact scores. [^9] Xie et al. (2024) use TF-IDF to rank the words. We find that the simpler approach of averaging the scores is sufficient for our purposes.
  4. 4.
    For each class, we select the 100 words with the highest impact scores for the analysis.
5 Results
Following the LOO method, we produce a list of 100 words with the highest impact scores per class in all settings. A manual inspection of the lists quickly shows a prevalence of place names and related items like Zürich or Österreicher (G ‘Austrian’). Therefore, as a first step, we count the amount of these words among the top 100 and find that, on average, 14% of words are local references. In terms of classification, these results are expected (see, e.g., Nasar et al., 2022) and it is apparent how these words are indicative of location given what they denote. Although it does not take a forensic linguist to understand the connection of these items to location, it is noteworthy that the classification models pick up on these words and that they match the region they are impactful for.
Given that the results are similar for classes based on country borders across the three settings, we focus our analysis on the 5-class setting. As a reminder, in this setting, Austria and Switzerland form individual classes, whereas Germany is split into three regions (northern, south-eastern and south-western). Generally, the results show that a large proportion of extracted words are regionalized and some are prototypical dialect items.
For Switzerland, we find both Swiss German and French lexical items to be the most impactful, such as Kei (SWG ‘not a/no’), Isch (SWG ‘is’) and bim (SWG ‘at the’), or pourquoi (FR ‘why’), Avec (FR ‘with’) and raison (FR ‘reason’). Whereas for Austria, the data shows that some of the most impactful lexical items are Oasch (AG ‘ass’), Gspusis (AG ‘affairs’), and Matura (AG ‘high-school (diploma)’). These three items are examples of a textualization of regional pronunciation, a regional item and a standard Austrian variant, respectively. Also, items that are relevant to the Jodel demographic, such as Jus (AG ‘law (studies)’), are extracted. This indicates that an automated classification between the three countries, for instance to clarify jurisdiction, seems reasonable.
For the German classes we find that a large proportion of top words appear textualized in standard German as opposed to more colloquial spellings including abbreviations and ellipses, which we may expect. Examples of this include Dankeschön (G ‘thank you’) or Vorname (G ‘first name’). Considering the division within Germany, we find that several forms of the verb gucken (G ‘to look/watch’) are impactful for the northern class, which is a variant we know to be regionalized and appearing in varieties in central and northern Germany (König, 2004, p. 235). For the south-east class the data shows items like Ritter (G ‘knight’), #traudel (female first name) and local beer types. For the south-west we find that items identified as relevant by Purschke and Hovy (2019), are also extracted by the LOO model, like Möppes (G ‘breasts’ or ‘female user’) and Lörres (G ‘penis’ or ‘male user’), although the authors argue that these forms are more Jodel- than region-specific. These items are also impactful for the Germany-class in the 3-class setting.
6 Conclusion
In this paper, we employed Xie et al. (2024)’s approach as a means of geolinguistic profiling to understand how a method like this could be applied in a forensic context given its explainability. We have found that the dialect classifiers outperform the random baseline by a large margin in all settings, but that accuracy decreases for settings with more closely-related classes. While we recognize that in forensic contexts the focus needs to be on false predictions and hard-to-classify cases, this paper considers the explainability of the approach. To this end we have found that the LOO model extracts meaningful regional features reflecting the variety used in the corresponding area. On average, 14% of extracted features are place names or similar items. While an analysis for a place name like Wien does not need a forensic linguist, extracting features that are not based on a linguist’s expertise is a valuable contribution of this approach even if it is not directly used for automated classification.
Limitations
For this paper we have worked with the geolocation of the Jodel posts as the gold label for the dialect regions used as classes in the classification task. However, there is noise in this data as people move and use varieties from different regions in the same place. Further analysis of the incorrect classifications may allow us to identify these instances.
For further work it may be beneficial to remove non-German varieties before training. Additionally, given the high percentage of place names and related lexical items, preprocessing to remove named entities (see, e.g., Darji et al., 2023) may help focus the extraction on dialectal lexical items.
Acknowledgements
Dana Roemling was supported by the UKRI ESRC Midlands Graduate School Doctoral Training Partnership ES/P000711/1.
We thank Dirk Hovy and Christoph Purschke for sharing their data with us.
We wish to acknowledge CSC – IT Center for Science, Finland, for computational resources.